Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2015-10-05

Observers: Joe Lyman, Heather Campbell
TAT: Joe Anderson, Cosimo Inserra
Data reducers: Morgan Fraser

EFOSC used for whole night. Wind constant at ~9-10m/s with persistent partial clouds that caused a couple of closures. Seeing quite erratic.

UT      Target                  Comment
========================================================================
23:20   LTT7379                 Spec standard
23:40                           Somewhat cloudy start, clearing during first target
23:55   PS15cem                 Classification (very young II!)
23:55                           Seeing ~1.3"
00:05   PS15cem                 Immediate follow up gr11 as per group's instructions
01:15                           Cirrus appeared
01:15   PSNJ23002463+0137368    Started followup but difficult to detect even in extended acq image, didn't attempt spectrum
01:20   PS15cez                 Classification - (Ia, post max)
01:45                           Seeing ~1.9"
01:45   SN2014cx                Follow up gr13 (1.5" slit, second exposure affected by poor conditions)
03:25                           Cloudy, dome closed
05:15                           Dome opened
05:15   PSNJ23002463+0137368    Follow up gr11
06:25                           Seeing improved ~1.2"
06:25   OGLE15lg                Classification (Ia around max)
06:55   ASASSN-14ha             Followup photometry (V- and part of R-band completed)
08:00                           Cloudy, dome closed

Comments